A single ticket to a magic show costs m dollars, but there is a discount of 25% per ticket if a group buys 20 or more tickets. The expression 20m - 20(0.25m) describes the total cost of buying 20 tickets for a group.


1
Expert's answer
2021-10-06T17:11:30-0400

cost of one tickets = M

If a group buys 20 tickets

then the cost would be "=20\times M=20M"

but we are given that if a group buys 20 or more tickets then will get a discount of 25% per tickets

since we are calculating the cost of 20 tickets

so the discount on the 1st ticket =25% of m

"=\frac{25}{100} \times M\\=0.25M"

so the final cost of 20 tickets "=20\times 0.25M"

therefore the total cost of buying 20 tickets fir group "=20M - 20(0.25M)"